OpenStudy (anonymous):

What percent of 50 is 18? show work please

OpenStudy (anonymous):

set up a proportion \[x /100=18/100\]

OpenStudy (anonymous):

I meant 18/50

OpenStudy (anonymous):

you could cross multiply and solve for x,but I think that it's easier to see that 50*2=100, so 18*2=x

OpenStudy (anonymous):

so x=36 18 is 36 percent of 50